# Integrable sub-Riemannian geodesic flows on the special orthogonal group

### v2

@@ Abstract.
We analyse the geometry of the rolling distribution on the special orthogonal group and show that almost all right-invariant sub-Riemannian metrics defined on this distribution have completely integrable geodesic flows. Our argument is an adaptation of the method used to establish integrability of the Riemannian metric arising from the $n$-dimensional rigid body: namely, by exhibiting a Lax pair and bi-Hamiltonian structure for the reduced equations of motion.
@@ 1. Introduction
The geodesic flow on a Riemannian manifold $M$ is generated by a Hamiltonian system on the cotangent bundle. A Riemannian metric is called integrable if the corresponding Hamiltonian system is integrable in the sense of Liouville; that is, if the the flow admits $n=\dim M$ independent integrals of motion which mutually commute with respect to the Poisson bracket on $T^{*}M$. In principle, one can then find the geodesics explicitly through quadrature. The literature surrounding integrable Riemannian metrics is vast (see for instance [6, 33, 29, 11, 10]) whereas less has been said about sub-Riemannian metrics, with much of the work focussed on Carnot groups [22, 25, 12]. A sub-Riemannian metric also defines a Hamiltonian system on the cotangent bundle, and so one can equally ask if this flow is integrable.
A famous example of an integrable Riemannian metric on $\mathbf{SO}(n)$ is that which arises from the kinetic energy of an $n$-dimensional rigid body. We will consider a related class of sub-Riemannian metrics defined on the right-invariant ‘rolling’ distribution $\mathcal{D}$ of tangent vectors to $\mathbf{SO}(n)$ of the form $\dot{g}=\Omega g$ where
@@
We prove this theorem by adapting the same methods which led Manakov to establish the complete integrability of the rigid body [23]. In particular, we consider the symplectic reduced space $\mathfrak{so}(n)^{*}$ and show that the corresponding reduced equations of motion admit a Lax pair formulation and a bi-Hamiltonian structure. This yields a collection of first integrals which can then be shown to define an integrable system thanks to the work of Mishchenko & Fomenko in [24].
Any sub-Riemannian problem may be reinterpreted as a kinematic non-holonomic optimal control problem. From this point of view, our problem is a special case of a more general class of problems defined on symmetric spaces considered in [7, 4]. There is also a wider literature concerning the related problem of rolling manifolds along each other, see for instance [19, 17]. In particular, we note that Theorem 1 was proven in the symmetric case where $A_{2}=\dots=A_{n}$ for $n=3$ in [3] and more generally in [20] where the sub-Riemannian problem corresponds to a ball rolling on a plane.
@@ 2.1. The Hamiltonian setting
on $T^{*}M$ is the Hamiltonian for the sub-Riemannian metric. It is known that the integral curves of the Hamiltonian flow when projected down to $M$ are, for sufficiently small arcs, the unique length-minimising admissible curves between their endpoints [26].
@@
for the left and right actions are given by right and left translation of covectors back to the identity, respectively. For left-invariant Hamiltonians the momentum $\mathbf{J}_{L}$ is constant along the Hamiltonian flow, and $\mathbf{J}_{R}$ serves as an orbit-quotient; and vice versa with the words left and right interchanged. The orbit space $\mathfrak{g}_{\pm}^{*}$ is Poisson with the bracket
(3) $\{f,g\}(M)=\pm\langle M,[df,dg]\rangle$
@@ 2.2. The rigid body
where $\mathbb{J}\colon\mathfrak{so}(n)\rightarrow\mathfrak{so}(n)^{*}$ is the inertia operator [30]. If we identify $\mathfrak{so}(n)$ with its dual via the trace pairing $\langle M,\Omega\rangle=\operatorname{Tr}(M^{T}\Omega)$ then
@@
where $\lambda$ is an arbitrary parameter. This equation implies that solution curves remain inside the same conjugacy class, and so $\operatorname{Tr}(M+\lambda J^{2})^{k}$ is constant for any power $k$. Since $\lambda$ was arbitrary it follows that each of the coefficients of $\lambda$ in the expansion
@@ 2.3. The sub-Riemannian metric
Left/right multiplication pushes $\mathfrak{p}$ forward to define a non-integrable left/right-invariant distribution on $\mathbf{SO}(n)$ equipped with a sub-Riemannian metric. The resulting length functional on the left-invariant distribution is unphysical in the sense that it is not the Lagrangian for a constrained physical body in $\mathbf{R}^{n}$. However, for the right-invariant distribution we can interpret the length functional as describing the length traced by the path of an $(n-1)$-dimensional ball rolling without slipping on $\mathbf{R}^{n-1}$ with metric
$ds^{2}=A_{2}dx_{2}^{2}+\dots+A_{n}dx_{n}^{2}.$
For this reason we shall work exclusively with the right-invariant distribution which we denote by $\mathcal{D}$ and refer to as the rolling distribution. We will later solve for the rolling motion in the case $n=3$.
@@
Notice that the sub-Riemannian Hamiltonian in (6) is the limit of the rigid-body Hamiltonian in (4) when $I_{kl}\rightarrow\infty$ for $k\neq 1$ and $A_{k}=I_{1k}$. Indeed, any sub-Riemannian metric on a distribution $\mathcal{D}$ can be interpreted as a limit of Riemannian penalty metrics in which velocity vectors not belonging to $\mathcal{D}$ become increasingly expensive. We note however that this limit is not physical since $2J_{1}=I_{12}+I_{13}-I_{23}$ necessarily limits to $\infty$- ∞ which contradicts the positive-definiteness of the mass matrix for realistic bodies. On the other hand, the metric on the left-invariant distribution does correspond to the optimal control of an underactuated $n$-dimensional rigid body, with control torques (actuators) about $(n-1)$ of the $n$ principal axes. For the underactuated rigid body problem see e.g. [7] and [4] and references therein.
@@ 2.4. Rolling a ball across a table
Place a ball on a table. The configuration of the ball is determined by its point of contact $(y,z)\in\mathbf{R}^{2}$ and a matrix $g\in\mathbf{SO}(3)$ whose columns are the basis vectors of an orthonormal frame fixed inside the ball. How can we roll the ball from one orientation $g_{0}$ to another $g_{1}$ in such a way that the curve of contact which the ball traces on the table has the least length with respect to the metric
@@
The rolling constraint implies that admissible tangent vectors to $\mathbf{SO}(3)$ must be of the form
(7) $\frac{dg}{dt}=\begin{pmatrix}0&-\Omega_{12}&\Omega_{13}\\ \Omega_{12}&0&0\\ -\Omega_{13}&0&0\end{pmatrix}g=\widehat{\Omega}g.$
This notation is chosen to coincide with the ‘hat-map’ whereby $\widehat{\Omega}v=\Omega\times v$ for $\Omega=(0,\Omega_{13},\Omega_{12})^{T}$. For such a motion, the path traced on the plane satisfies
(8) $\frac{d}{dt}\begin{pmatrix}y\\ z\end{pmatrix}=\begin{pmatrix}\Omega_{12}\\ -\Omega_{13}\end{pmatrix},$
where we are supposing for simplicity that the ball has radius $\pi^{-1}$. For an admissible curve $g(t)$ the length of the path traced by the ball is therefore
[…]
@@
$H_{\text{sR}}(M)=\frac{1}{2}\left(\frac{M_{12}^{2}}{A_{2}}+\frac{M_{13}^{2}}{A_{3}}\right)$
for $M=(M_{23},M_{13},M_{12})^{T}$. After reduction this descends to a Lie-Poisson system in $\mathfrak{so}(3)^{*}_{+}\cong\mathbf{R}^{3}$ with equation of motion
@@
The Hamiltonian $H_{\text{sR}}$ and Casimir function $|M|^{2}$ are both constants of motion. The intersections of their levels sets therefore give the solution curves: elliptic cylinders for $H_{\text{sR}}$ constant, and spheres of constant $|M|^{2}$, as shown on the right-hand side in Figure 1.
For the initial conditions $M_{12}(0)=\sqrt{C-A_{2}}$, $M_{12}(0)=\sqrt{A_{2}}$ and $M_{13}(0)=0$ we can solve (10) exactly in terms of Jacobi elliptic functions, finding
(11) $M_{12}(t)=\sqrt{A_{2}}\operatorname{cn}(u,k)\quad\text{and}\quad M_{13}(t)=\sqrt{A_{3}}\operatorname{sn}(u,k),$
@@
The tangent vector $\Omega$ is related to the covector $M$ via $M_{12}=A_{2}\Omega_{12}$ and $M_{13}=A_{2}\Omega_{13}$. We can use this relation together with (8) to find the path that the ball traces by integrating (11) (with the aid of [13]) to find that
@@
The rolling problem we have discussed is related to an interview question asked in the 1950s at the University of Oxford [16], first solved in [2] (see also [18]). Given two points $p_{0}$ and $p_{1}$ on the table, and two configurations $g_{0}$ and $g_{1}$ of the ball, what is the shortest length path which rolls $(p_{0},g_{1})$ to $(p_{1},g_{1})$? Our problem differs from this in two respects. Firstly, we do not care about connecting two points on the table, only the two configurations of the ball; and secondly, the metric on the table is not the standard Euclidean metric. It is curious to compare the curves in Eq. (12) with those for the Euclidean case $A_{2}=A_{3}$ where the ball is found to travel across the table in straight lines and circular arcs.
@@ 3.1. Manakov integrals in the limit
Following on from Remark 2.1 we may interpret the sub-Riemannian metric as a limit of Riemannian metrics for the rigid body. Since the rigid body is known to be integrable for distinct principle moments of inertia, does it follow that its limit is also integrable? Our initial heuristic for finding first integrals was to inspect the Manakov integrals in such a limit. However, the integrals are found to diverge or become dependent. To overcome this, we now show how to define a collection of integrals which are functionally equivalent to the Manakov integrals, but which survive in the limit. We will later show that these are in fact first integrals of the sub-Riemannian flow.
To define this limit more precisely consider a rigid-body parametrised by $s$ whose mass matrix $J^{s}$ is diagonal with entries
@@
In this way $I_{1k}=A_{k}$ for all $s$, and the $I_{kl}\rightarrow\infty$ as $s\rightarrow\infty$ for $k\neq 2$. The Hamiltonian therefore limits to the sub-Riemannian Hamiltonian $H_{\text{sR}}$. Strictly speaking this metric does not represent a physical body since $J^{s}_{1}$ is negative, however this does not affect the definition of the Manakov integrals. We may rewrite
@@
for arbitrary $\sigma$. In the next section we will verify that these are indeed first integrals of the sub-Riemannian geodesic flow.
@@ 3.2. A Lax pair and bi-Hamiltonian structure
(13) $\frac{d}{dt}M=[dH_{\text{sR}},M],$
in $\mathfrak{so}_{+}(n)^{*}$ where the derivative $dH_{\text{sR}}$ is understood to be an element of the Lie algebra. It will be helpful to decompose
$M=\begin{pmatrix}0&-\mu^{T}\\ \mu&\hat{M}\end{pmatrix}$
into the $(n-1)$-dimensional column vector $\mu$ and the skew-symmetric matrix $\hat{M}$ belonging to $\mathfrak{so}(n-1)$. In a slight abuse of notation, by letting $\hat{A}^{-1}$ denote the matrix $\text{diag}(A_{2}^{-1},\dots,A_{n}^{-1})$ the Hamiltonian is conveniently written
@@
(14) $dH_{\text{sR}}=\begin{pmatrix}0&-\mu^{T}\hat{A}^{-1}\\ \hat{A}^{-1}\mu&0\end{pmatrix}.$
@@
$\frac{d}{dt}\mu$ $=-\hat{M}\hat{A}^{-1}\mu,$
$\frac{d}{dt}\hat{M}$ $=[\mu\mu^{T},\hat{A}^{-1}].$
This admits a Lax pair formulation,
(15) $\frac{d}{dt}(M+\sigma A)=[dH_{\text{sR}}+\sigma B,M+\sigma A],$
for $\sigma$ an arbitrary parameter and $B=\text{diag}(0,1,\dots,1)$. It follows that $M+\sigma A$ evolves within the same conjugacy class. The traces of $(M+\sigma A)^{k}$ are therefore integrals of motion. Since $\sigma$ is arbitrary, the coefficients of $\sigma$ in the expansion
@@
We now turn our attention to establishing a bi-Hamiltonian structure for the equations of motion. For any symmetric matrix $S$ there is an alternative Lie bracket on $\mathfrak{so}(n)$ given by
$[X,Y]_{S}=XSY-YSX.$
We can use this to modify (3) to produce an alternative Poisson structure on $\mathfrak{so}(n)^{*}$ given by
$\{f,g\}_{S}=\langle M,[df,dg]_{S}\rangle.$
@@
holds for all $\varphi$. The Hamiltonian system generated by $H$ is therefore equivalent to the system generated by $F$ with respect to the altered Poisson structure. If $S$ is not a scalar multiple of the identity then we say that this system is bi-Hamiltonian.
@@
(17) $F(M)=\operatorname{Tr}\left(\mu^{T}\hat{A}^{-1}\mu\right)+\frac{1}{2}\operatorname{Tr}\left(\hat{M}^{T}\hat{A}^{-1}\hat{M}\hat{A}^{-1}\right)$
with respect to the altered Poisson structure {,}$\{\leavevmode\nobreak\ ,\leavevmode\nobreak\ \}_{A}$.
@@
(18) $\langle M,[X,dH_{\text{sR}}]\rangle=\langle M,[X,dF]_{A}\rangle$
@@
$\langle M,[X,dH_{\text{sR}}]\rangle=\langle X,[dH_{\text{sR}},M],\rangle\text{and }\langle M,[X,dF]_{A}\rangle=\langle X,[A,M]_{dF}\rangle.$
@@
(19) $\langle X,[dH_{\text{sR}},M]\rangle=\langle X,[A,M]_{dF}\rangle.$
@@
(20) $dF=\begin{pmatrix}0&-\mu^{T}\hat{A}^{-2}\\ \hat{A}^{-2}\mu&\hat{A}^{-1}\hat{M}\hat{A}^{-1}\end{pmatrix}.$
which when combined with (14) allows us to compute and verify the important identity $[dH_{\text{sR}},M]=[A,M]_{dF}$, from which (19) follows.

### v3

@@ Abstract.
We analyse the geometry of the rubber-rolling distribution on the special orthogonal group and show that almost all the normal geodesics of any right-invariant sub-Riemannian metric defined on this distribution are completely integrable. Our argument is an adaptation of the method used to establish integrability of the Riemannian metric arising from the $n$-dimensional rigid body: namely, by exhibiting a Lax pair and bi-Hamiltonian structure for the reduced equations of motion.
@@ 1. Introduction
The geodesic flow on a Riemannian manifold $M$ is generated by a Hamiltonian system on the cotangent bundle. A Riemannian metric is called integrable if the corresponding Hamiltonian system is integrable in the sense of Liouville, that is, if the flow admits $n=\dim M$ independent integrals of motion which mutually commute with respect to the Poisson bracket on $T^{*}M$. In principle, one can then find the geodesics explicitly through quadrature. The literature surrounding integrable Riemannian metrics is vast (see, for instance, [8, 40, 35, 13, 12]) whereas less has been said about sub-Riemannian metrics, with much of the work focussed on Carnot groups [28, 31, 14]. A sub-Riemannian metric also defines a Hamiltonian system on the cotangent bundle, and so one can equally ask if this flow is integrable.
A famous example of an integrable Riemannian metric on $\mathbf{SO}(n)$ is that which arises from the kinetic energy of an $n$-dimensional rigid body. We will consider a related class of sub-Riemannian metrics defined on the right-invariant ‘rubber-rolling’ distribution $\mathcal{D}$ of tangent vectors to $\mathbf{SO}(n)$ of the form $\dot{g}=\Omega g$ where
@@
We prove this theorem by adapting the same methods which led Manakov to establish the complete integrability of the rigid body [29]. In particular, we consider the symplectic reduced space $\mathfrak{so}(n)^{*}$ and show that the corresponding reduced equations of motion admit a Lax pair formulation and a bi-Hamiltonian structure. This yields a collection of first integrals which can then be shown to define an integrable system thanks to the work of Mishchenko and Fomenko in [30]. The idea of applying Mishchenko and Fomenko’s arguments to a sub-Riemannian metric instead of an ordinary Riemannian metric was pointed out in [9] and [16].
Any sub-Riemannian problem may be reinterpreted as a kinematic non-holonomic optimal control problem. From this point of view, our problem is a special case of a more general class of problems defined on symmetric spaces considered in [9, 6]. There is also a wider literature concerning the related problem of rolling manifolds along each other, see for instance [24, 20]. In particular, we note that Theorem 1 was proven in the symmetric case where $A_{2}=\dots=A_{n}$ for $n=3$ in [4] and more generally in [25] where the sub-Riemannian problem corresponds to a ball rolling on a plane without spinning.
@@ 2.1. The Hamiltonian setting
on $T^{*}M$ is the Hamiltonian for the sub-Riemannian metric. It is known that the integral curves of the Hamiltonian flow when projected down to $M$ are, for sufficiently small arcs, the unique length-minimising admissible curves between their endpoints.
@@
for the left and right actions are given by right and left translation of covectors back to the identity, respectively. For left-invariant Hamiltonians, the momentum $\mathbf{J}_{L}$ is constant along the Hamiltonian flow, and $\mathbf{J}_{R}$ serves as an orbit-quotient; and vice versa with the words left and right interchanged. The orbit space $\mathfrak{g}_{\pm}^{*}$ is Poisson with the bracket
(3) $\{f,g\}(M)=\pm\langle M,[df|_{M},dg_{M}]\rangle$
@@
For a more detailed background in sub-Riemannian geometry, we recommend the references [32], [2], and [6], and [34] for those aspects concerning Poisson geometry.
@@ 2.2. The rigid body
where $\mathbb{J}\colon\mathfrak{so}(n)\rightarrow\mathfrak{so}(n)^{*}$ is the inertia operator [36]. If we identify $\mathfrak{so}(n)$ with its dual via the trace pairing
$\langle M,\Omega\rangle=\frac{1}{2}\operatorname{Tr}(M^{T}\Omega)$
then
@@
where $\lambda$ is an arbitrary parameter. This equation, found originally by Manakov [29], implies that solution curves remain inside the same conjugacy class, and so $\operatorname{Tr}(M+\lambda J^{2})^{k}$ is constant for any power $k$. Since $\lambda$ was arbitrary it follows that each of the coefficients of $\lambda$ in the expansion
@@ 2.3. The sub-Riemannian metric
Left/right multiplication pushes $\mathfrak{p}$ forward to define a non-integrable left/right-invariant distribution on $\mathbf{SO}(n)$ equipped with a sub-Riemannian metric. The resulting length functional on the left-invariant distribution corresponds to the $n$-dimensional Suslov problem which concerns the motion of a rigid body with constrained angular momentum. However, the dynamics of this system are non-holonomic, and as explained in [42] do not correspond to a system of Lie-Poisson equations on $\mathfrak{so}(n)_{-}^{*}$ but instead to a system of so-called Euler-Poincaré-Suslov equations. Therefore, we focus our attention instead on the right-invariant distribution whose length functional may be interpreted as describing the length traced by the path of an (n1)(n-1)-dimensional ball which rolls without sliding or spinning on $\mathbf{R}^{n-1}$ with metric
(6) $ds^{2}=A_{2}dy_{2}^{2}+\dots+A_{n}dy_{n}^{2}.$
We denote this right-invariant distribution by $\mathcal{D}$ and refer to it as the rubber-rolling distribution, a term we have taken from [17] and [26], so called since the ball is not allowed to spin about its point of contact. We will later solve for the rolling motion in the case $n=3$.
@@
Notice that the sub-Riemannian Hamiltonian in (7) is the limit of the rigid-body Hamiltonian in (4) when $I_{kl}\rightarrow\infty$ for $k\neq 1$ and $A_{k}=I_{1k}$. Indeed, any sub-Riemannian metric on a distribution $\mathcal{D}$ can be interpreted as a limit of Riemannian penalty metrics in which velocity vectors not belonging to $\mathcal{D}$ become increasingly expensive. We note, however, that this limit is not physical since $2J_{1}=I_{12}+I_{13}-I_{23}$ necessarily limits to -$\infty$, which contradicts the positive-definiteness of the mass matrix for realistic bodies. On the other hand, the metric on the left-invariant distribution does correspond to the optimal control of an underactuated $n$-dimensional rigid body, with control torques (actuators) about (n1)(n-1) of the $n$ principal axes. For the underactuated rigid body problem, see e.g. [9] and [6] and references therein.
@@ 2.4. Rolling a ball across a table
Place a rubber ball on a table. The configuration of the ball is determined by its point of contact $(y,z)\in\mathbf{R}^{2}$ and a matrix $g\in\mathbf{SO}(3)$ whose columns are the basis vectors of an orthonormal frame fixed inside the ball. How can we roll the ball from one orientation $g_{0}$ to another $g_{1}$ in such a way that the curve of contact which the ball traces on the table has the least length with respect to the metric
@@
The no-spinning constraint implies that admissible tangent vectors to $\mathbf{SO}(3)$ must be of the form
(8) $\frac{dg}{dt}=\begin{pmatrix}0&-\Omega_{z}&\Omega_{y}\\ \Omega_{z}&0&0\\ -\Omega_{y}&0&0\end{pmatrix}g=\widehat{\Omega}g.$
This notation is chosen to coincide with the ‘hat-map’ whereby $\widehat{\Omega}v=\Omega\times v$ for $\Omega=(0,\Omega_{y},\Omega_{z})^{T}$. For such a motion, the path traced on the plane satisfies
(9) $\frac{d}{dt}\begin{pmatrix}y\\ z\end{pmatrix}=\begin{pmatrix}\Omega_{z}\\ -\Omega_{y}\end{pmatrix}$,
where we are supposing for simplicity that the ball has a unit radius. For an admissible curve $g(t)$ the length of the path traced by the ball is therefore
(10) $\int\sqrt{A_{2}|\Omega_{z}|^{2}+A_{3}|\Omega_{y}|^{2}}\leavevmode\nobreak\ dt.$
@@
$H_{\text{sR}}(M)=\frac{1}{2}\left(\frac{M_{z}^{2}}{A_{2}}+\frac{M_{y}^{2}}{A_{3}}\right)$
for $M=(M_{x},M_{y},M_{z})^{T}$. After reduction this descends to a Lie-Poisson system in $\mathfrak{so}(3)^{*}_{+}\cong\mathbf{R}^{3}$ with equation of motion
@@
The Hamiltonian $H_{\text{sR}}$ and Casimir function $|M|^{2}$ are both constants of motion. The intersections of their levels sets therefore give the solution curves: elliptic cylinders for $H_{\text{sR}}$ constant, and spheres of constant $C=|M|^{2}$, as shown on the right-hand side in Figure 1.
For the initial conditions $M_{x}(0)=\sqrt{C-A_{2}}$, $M_{y}(0)=0$ , and $M_{z}(0)=\sqrt{A_{2}}$ we can solve (11) exactly in terms of Jacobi elliptic functions, finding
(12) $M_{y}(t)=\sqrt{A_{3}}\operatorname{sn}(u,k)\quad\text{and}\quad M_{z}(t)=\sqrt{A_{2}}\operatorname{cn}(u,k),$
@@
The tangent vector $\Omega$ is related to the covector $M$ via $M_{y}=A_{3}\Omega_{y}$ and $M_{z}=A_{2}\Omega_{z}$. We can use this relation together with (9) to find the path that the ball traces by integrating (12) (with the aid of [15]) to find that
@@
The rolling problem we have discussed is related to an interview question asked in the 1950s at the University of Oxford [19], first solved in [3] (see also [22] and [23]). Given two points $p_{0}$ and $p_{1}$ on the table, and two configurations $g_{0}$ and $g_{1}$ of the ball, what is the shortest length path which rolls $(p_{0},g_{1})$ to $(p_{1},g_{1})$? Our problem differs from this in two respects. Firstly, we are not concerned with connecting two points on the table, but rather with the two configurations of the ball; and secondly, the metric on the table is not the standard Euclidean metric. It is curious to compare the curves in Eq. (13) with those for the Euclidean case $A_{2}=A_{3}$ where the ball is found to travel across the table in straight lines and circular arcs.
We conclude by noting that we have only solved for the path which the ball traces on the table. In order to find the explicit sub-Riemannian geodesics on the group $\mathbf{SO}(3)$ Eq. (8) must be integrated. These geodesics have been found explicitly in [5].
@@ 3.1. Manakov integrals in the limit
Following on from Remark 2.1 we may interpret the sub-Riemannian metric as a limit of Riemannian metrics for the rigid body. Since the rigid body is known to be integrable for distinct principal moments of inertia, does it follow that its limit is also integrable? Our initial heuristic for finding first integrals was to inspect the Manakov integrals in such a limit. However, the integrals are found to diverge or become dependent. To overcome this, we now show how to define a collection of integrals which are functionally equivalent to the Manakov integrals, but which survive in the limit. We will later show that these are in fact first integrals of the sub-Riemannian flow.
To define this limit more precisely, consider a rigid-body parametrised by $s$ whose mass matrix $J^{s}$ is diagonal with entries
@@
In this way $I_{1k}=A_{k}$ for all $s$, and the $I_{kl}\rightarrow\infty$ as $s\rightarrow\infty$ for $k\neq 1$. The Hamiltonian therefore limits to the sub-Riemannian Hamiltonian $H_{\text{sR}}$. Strictly speaking, this metric does not represent a physical body since $J^{s}_{1}$ is negative, however this does not affect the definition of the Manakov integrals. We may rewrite
@@
for arbitrary $\sigma$. In the next section, we will verify that these are indeed first integrals of the sub-Riemannian geodesic flow.
@@ 3.2. A Lax pair and bi-Hamiltonian structure
(14) $\frac{d}{dt}M=[dH_{\text{sR}}|_{M},M],$
in $\mathfrak{so}_{+}(n)^{*}$ where the derivative $dH_{\text{sR}}$ evaluated at $M$ is understood to be an element of the Lie algebra. It will be helpful to decompose
(15) $M=\begin{pmatrix}0&-\mu^{T}\\ \mu&\widehat{m}\end{pmatrix}$
into the (n1)(n-1)-dimensional column vector $\mu$ and the skew-symmetric matrix $\widehat{m}$ belonging to $\mathfrak{so}(n-1)$. In a slight abuse of notation, by letting $\hat{A}^{-1}$ denote the matrix $\text{diag}(A_{2}^{-1},\dots,A_{n}^{-1})$ the Hamiltonian is conveniently written
@@
(16) $dH_{\text{sR}}|_{M}=\begin{pmatrix}0&-\mu^{T}\hat{A}^{-1}\\ \hat{A}^{-1}\mu&0\end{pmatrix}$.
@@
(17) $\begin{split}\frac{d}{dt}\mu&=-\widehat{m}\hat{A}^{-1}\mu,\\ \frac{d}{dt}\widehat{m}&=[\mu\mu^{T},\hat{A}^{-1}].\end{split}$
Proposition 3.1.
The equations of motion admit a Lax pair formulation,
(18) $\frac{d}{dt}(M+\sigma A)=[dH_{\textup{sR}}|_{M}+\sigma B,M+\sigma A],$
for $\sigma$ an arbitrary parameter, $A=\textup{diag}(0,A_{2},\dots,A_{n})$, and $B=\textup{diag}(0,1,\dots,1)$.
Proof.
This follows from the important identity
(19) $dH_{\textup{sR}}|_{M},A]=[M,B]$
which is verified by direct calculation. ∎
It follows that $M+\sigma A$ evolves within the same conjugacy class. The traces of $(M+\sigma A)^{k}$} are therefore integrals of motion. Since $\sigma$ is arbitrary, the coefficients of $\sigma$ in the expansion
@@
We now turn our attention to establishing a bi-Hamiltonian structure for the equations of motion. This can be used to give an elegant proof of the involutivity of the sub-Riemannian integrals by closely adapting the work of [33]. We will, however, prove this using an alternative method, owing to Mishchenko and Fomenko, in the next section. Nonetheless, we include the work below since it constitutes an interesting example of a sub-Riemannian flow admitting a bi-Hamiltonian structure.
We begin by observing that for any matrix $S$ there is an alternative Lie bracket on the space of matrices given by
$X,Y]_{S}=XSY-YSX.$
If $S$ is symmetric then this restricts to a Lie bracket on $\mathfrak{so}(n)$ which we can use to modify (3) to produce an alternative Poisson structure on $\mathfrak{so}(n)^{*}$ given by
$\{f,g\}_{S}=\langle M,[df|_{M},dg|_{M}]_{S}\rangle.$
@@
holds for all $\varphi$. The Hamiltonian system generated by $H$ is therefore equivalent to the system generated by $F$ with respect to the altered Poisson structure. If $S$ is not a scalar multiple of the identity, then we say that this system is bi-Hamiltonian.
@@
(21) $F(M)=\operatorname{Tr}\left(\mu^{T}\hat{A}^{-1}\mu\right)+\frac{1}{2}\operatorname{Tr}\left(\widehat{m}^{T}\hat{A}^{-1}\widehat{m}\hat{A}^{-1}\right)$
with respect to the altered Poisson structure $\{\leavevmode\nobreak\ ,\leavevmode\nobreak\ \}_{A}$ where $A=\textup{diag}(0,A_{2},\dots,A_{n})$.
@@
(22) $\langle M,[X,dH_{\text{sR}}|_{M}]\rangle=\langle M,[X,dF|_{M}]_{A}\rangle$
@@
$\langle M,[X,dH_{\text{sR}}|_{M}]\rangle=\langle X,[dH_{\text{sR}}|_{M},M],\rangle\text{ and }\langle M,[X,dF|_{M}]_{A}\rangle=\langle X,[A,M]_{dF|_{M}}\rangle.$
@@
(23) $\langle X,[dH_{\text{sR}}|_{M},M]\rangle=\langle X,[A,M]_{dF|_{M}}\rangle.$
@@
(24) $dF|_{M}=\begin{pmatrix}0&-\mu^{T}\hat{A}^{-2}\\ \hat{A}^{-2}\mu&\hat{A}^{-1}\widehat{m}\hat{A}^{-1}\end{pmatrix}$
which when combined with (16) allows us to compute and verify the important identity $dH_{\text{sR}}|_{M},M]=[A,M]_{dF|_{M}}$ from which (23) follows.